OpenStudy (anonymous):

Consider the equations y=1x√ and y=x−2. The approximate solution of this system of equations is

OpenStudy (anonymous):

do you know how to do this one?? @peachpi

OpenStudy (anonymous):

what is the first equation? use the equation tool below

OpenStudy (anonymous):

\[y=1/\sqrt{x}\]

OpenStudy (anonymous):

\[y=\frac{1}{\sqrt{x}}\\ y=x-2\] solve \[\frac{1}{\sqrt{x}}=x-2\]

OpenStudy (anonymous):

that is where i get stuck

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Multiply both sides by √x, then I think it might be easier to change the variable and solve the resulting polynomial equation